# Enhancing online estimation of CBC parameters with the low-latency MBTA analysis

### v1

@@ 2.2 Fast template placement
  • projecting this bank using the metric of the template that maximizes the posterior probability (Equation 9).
@@ 3.1 MBTA detector-frame PE
Under the hypothesis that a CBC signal is present in the data $d$, the probability that the source matches the parameters $\boldsymbol{\lambda}$ of a given template $T$ is given by Bayes theorem, as the product of the likelihood and the template prior probability:
$\mathrm{p}(T(\boldsymbol{\lambda})\mid d)\propto\mathrm{p}(d\mid T(\boldsymbol{\lambda}))\cdot\mathrm{p}(T(\boldsymbol{\lambda})).$ (9)
@@
$\mathrm{p}(d\mid T({\boldsymbol{\lambda}_{\mathrm{int}}}))\propto\exp\left(-\sum\limits_{i}\mathrm{SNR}_{\mathrm{opt},i}^{2}\left(1-\frac{\mathrm{SNR}_{T,i}}{\mathrm{SNR}_{\mathrm{opt},i}}\right)\right),$ (10)
@@
The second term of Equation 9 is the prior probability density function (PDF) of the source parameters. To ensure consistency with the parameters reported by the collaboration in [6], we adopt the same agnostic priors: uniform distributions in detector-frame individual masses and spin magnitudes, with isotropic spin orientations.
Since our samples are drawn from the SNR optimizer bank, they follow a specific distribution in chirp times. Therefore, the reweighting procedure described in Appendix B is applied. Figure 2 highlights the differences between the template distribution and their prior probability, illustrating that the chosen prior favors symmetric masses and low effective spins, while the bank covers a broader parameter space. Figure 3a shows the detector-frame posterior probability evaluated for the parameters of each template filtered during the second iteration of the MBTA SNR optimizer, applied to the data associated with GW230529_181500 [2].
Figure 2: Prior probability of the templates used in the first iteration of the SNR optimizer, initiated by the MBTA trigger for GW230529_181500 [2], used alongside recovered SNRs to infer source parameters. The distributions are shown for the detector-frame chirp mass ($m_{\text{chirp}}^{d}$), the symmetric mass ratio ($\eta$), and the effective spin ($\chi_{\mathrm{eff}}$). The plots display the marginal 1D (diagonal) and 2D (off-diagonal) distributions. Black lines indicate the distributions of the templates used for filtering. Blue lines and dots show the same distributions reweighted according to Equation 19.
@@ 3.2 MBTA source-frame PE
The source properties can be inferred from mass and spin estimates given by Equation 9. However, because of the Universe’s expansion, the masses measured in the detector-frame appear larger due to redshift $z$ [43]. Since our PE method only probes the intrinsic parameters, it does not infer the source distance, nor the redshift. We therefore rely on the source localization provided in low-latency by the Bayestar algorithm [63], applied to the maximum likelihood template. Following the procedure in Appendix C, we convert the detector-frame mass and spin joint PDF into source-frame using the marginal luminosity distance PDF provided by Bayestar and assuming a standard cosmological model, described in [16].
@@
(b) Source-frame. Samples are obtained by rejection sampling of the calculated PDF.
@@ 3.3 Source classification and properties
The resulting source-frame PDF can be used to update the source classification and properties publicly provided to aid for inform EM surveys of the candidates.
Source classification is expressed as a set of probabilities representing the possible origin of the signal. Four mutually exclusive categories are considered: $p_{\text{terr}}$ for a non-astrophysical (terrestrial) origin, $p_{\text{BNS}}$ for both component masses between $1~\mathrm{M_{\odot}}$ and $3~\mathrm{M_{\odot}}$, $p_{\text{NSBH}}$ for a binary with one mass in this range and the other above $3~\mathrm{M_{\odot}}$, and $p_{\text{BBH}}$ for both components above $3~\mathrm{M_{\odot}}$ [46]. Since the MBTA SNR optimizer relies on a significant detection and does not re-estimate candidate significance, the value of $p_{\text{terr}}$ is taken from the online search [20, 19]. By integrating the source-frame joint individual mass PDF, marginalized over spin, the remaining probability is redistributed among the three astrophysical categories. Since standard LVK searches do not include a $p_{\text{astro}}$ sub-solar mass (SSM) category, the portion of the PDF corresponding to SSM components is excluded from this source classification process.
The collaboration also provides a set of probabilities inferring the possible properties of the source [46]. While the method for estimating these properties is shared across all LVK search pipelines [28], we show that the MBTA SNR optimizer PE process can be used to produce our own estimate. The HasSSM and HasMassGap probabilities can be computed by directly integrating the source-frame mass PDF over the regions where at least one component has a mass below $1~\mathrm{M_{\odot}}$, and between $3~\mathrm{M_{\odot}}$ and $5~\mathrm{M_{\odot}}$, respectively. The HasNS and HasRemnant definitions are dependent on neutron star models. Therefore, the limits for neutron star and remnant masses are marginalized over several neutron star equations of state [44] to weight the source-frame PDF before integration. In this process, neutron star spins in NSBH systems are fixed to zero, while black hole spins are allowed to vary according to the measured PDF.
@@ 4.3 Parameter recovery
For further validation, we applied our PE method to the O4a-events and compared the 2D posterior PDFs of the source-frame component masses obtained with our method to those provided by the LVK collaboration. We perform this comparison by evaluating the overlap $\mathcal{O}$ between the two binned distributions—denoted $p_{\mathrm{MBTA}}$ and $p_{\mathrm{LVK}}$—given by [25]:
@@ 4.5 Source classification and properties
Releasing the SNR optimizer classifications and properties would be appropriate even in the absence of an $\mathrm{SNR}$ gain, as they remain consistent with the results shown in Figures 10b and 11, the PE procedure—based on posterior PDF integration—being robust to small $\mathrm{SNR}$ variations.
@@ Acknowledgments
We express our gratitude to Gijs Nelemans and Jérôme Novak for their work in carrying out the internal LVK review of the MBTA pipeline and for their helpful comments and suggestions on the method presented in this paper. We thank our LVK collaborators from the CBC and low-latency groups for constructive comments. This analysis exploits the resources of the computing facility at the EGO-Virgo site, and of the Computing Center of the Institut National de Physique Nucléaire et Physique des Particules (CC-IN2P3 / CNRS). This research has made use of data or software obtained from the Gravitational Wave Open Science Center (gwosc.org), a service of the LIGO Scientific Collaboration, the Virgo Collaboration, and KAGRA. This material is based upon work supported by NSF’s LIGO Laboratory which is a major facility fully funded by the National Science Foundation, as well as the Science and Technology Facilities Council (STFC) of the United Kingdom, the Max-Planck-Society (MPS), and the State of Niedersachsen/Germany for support of the construction of Advanced LIGO and construction and operation of the GEO600 detector. Additional support for Advanced LIGO was provided by the Australian Research Council. Virgo is funded, through the European Gravitational Observatory (EGO), by the French Centre National de Recherche Scientifique (CNRS), the Italian Istituto Nazionale di Fisica Nucleare (INFN) and the Dutch Nikhef, with contributions by institutions from Belgium, Germany, Greece, Hungary, Ireland, Japan, Monaco, Poland, Portugal, Spain. KAGRA is supported by Ministry of Education, Culture, Sports, Science and Technology (MEXT), Japan Society for the Promotion of Science (JSPS) in Japan; National Research Foundation (NRF) and Ministry of Science and ICT (MSIT) in Korea; Academia Sinica (AS) and National Science and Technology Council (NSTC) in Taiwan.
@@ A.2 Universal bank generation
As discussed at the end of Section 2, we consider two different approaches. In the first iteration of the SNR optimizer, a constant minimal match of $98\%$ is used, resulting in templates uniformly distributed in dimensionless chirp times. At this stage, $4000$ potential templates are retained, represented in the top row of Figure 1. In the second iteration, templates are distributed to focus around the most probable point identified by the first pass. We adopt a sampling strategy designed to match the template distribution to the likelihood expressed in Equation 10. Templates are placed such that the probability of finding one at a given match $M$ with respect to the bank center, characterized by $\boldsymbol{\theta}_{O}$, is given by:
$p_{T}(\boldsymbol{\theta})=p_{0}\exp\!\left[-\tilde{\rho}^{2}\left(1-M(\boldsymbol{\theta}_{O},\boldsymbol{\theta})\right)\right],$ (16)
@@
with $\bar{\rho}$ denoting the mean SNR measured across detectors for the template yielding the highest network $\mathrm{SNR}$, $k$ a safety factor (typically set to $0.4$) used to artificially broaden the bank, and $\tilde{\rho}_{\max}$ a parameter meant to set an upper limit on $\tilde{\rho}$, fixed to $15$ to ensure that the second iteration bank encompasses more than one template from the first iteration. The normalization constant $p_{0}$ is chosen so that the resulting bank contains $7000$ potential templates.
@@ A.3 Sample projection
Appendix B Prior reweighting
In standard LVK parameter inference, an agnostic prior distribution is typically assumed. This prior is taken to be uniform within some plausible range in individual detector-frame component masses and spin magnitudes, while isotropic in spin orientations [5]. However, the parameter estimation technique presented here relies on samples distributed according to a predefined PDF $p_{T}({\boldsymbol{\theta}})$, where the parameters are expressed in terms of the dimensionless chirp-times $\boldsymbol{\theta}\equiv(\theta_{0},\theta_{3},\theta_{3s})$, defined in Equation 7. This distribution either is uniform in those parameters for the first SNR optimizer iteration, or follows Equation 16 for the second iteration. To achieve the desired prior distribution, we apply the reweighting described in [27]:
$p(T(\boldsymbol{\lambda_{\mathrm{int}}=\theta}))=\frac{p(\boldsymbol{\theta}\mid m_{1}^{d},m_{2}^{d},\chi_{1},\chi_{2},\phi_{1},\phi_{2})}{p_{T}({\boldsymbol{\theta}})},$ (19)
where the numerator represents the probability of drawing a given set of dimensionless chirp-times $\boldsymbol{\theta}$, conditioned on the prior probability of the individual detector-frame component masses $(m_{1}^{d},m_{2}^{d})$, spin magnitudes $(\chi_{1},\allowbreak\chi_{2})$, and spin orientation angles $(\phi_{1},\phi_{2})$ relative to the orbital angular momentum of the binary.
Since individual spins are constrained to be identical by the bank generation process, the spin dependence in $\boldsymbol{\theta}$ is encapsulated in a single spin component aligned with the orbital angular momentum:
$\chi^{z}=\chi_{1}\cos(\phi_{1})=\chi_{2}\cos(\phi_{2}).$ (20)
Then, the numerator in Equation 19 simplifies to:
$\begin{split}&p(\boldsymbol{\theta}\mid m_{1}^{d},m_{2}^{d},\chi_{1},\chi_{2},\phi_{1},\phi_{2})\\ &=p(\boldsymbol{\theta}\mid m_{1}^{d},m_{2}^{d},\chi^{z})\\ &=|J|\cdot p_{m_{1}^{d},m_{2}^{d},\chi^{z}}(m_{1}^{d},m_{2}^{d},\chi^{z}),\end{split}$ (21)
@@ Appendix B Prior reweighting
with $\omega_{\eta}^{2}=4\pi^{2}f_{0}^{2}\sqrt{1-4\eta}$. Assuming a standard prior where masses and spins are uncorrelated and masses are uniformly distributed, the prior simplifies:
$\begin{split}p_{m_{1}^{d},m_{2}^{d},\chi^{z}}(m_{1}^{d},m_{2}^{d},\chi^{z})\propto p_{\chi^{z}}(\chi^{z}).\end{split}$ (23)
@@
$p_{\chi^{z}}(\chi^{z})=-\frac{1}{2}\log_{e}\left(|\chi^{z}|\right),$ (24)
Combining Equation 22 and Equation 24 gives the numerator in Equation 19. Reweighting by the PDF used for template placement provides the prior PDF. To prevent the prior PDF from diverging at $\eta=0.25$ (Equation 22) or $\chi^{z}=0$ (Equation 24), the parameters used in its evaluation are clipped to $q=m_{2}/m_{1}\leq 0.99$ and $|\chi^{z}|\geq 0.001$. An illustration of such PDF is shown in Figure 2.
@@ Appendix C Detector-frame into source-frame posterior conversion
The parameter estimation method presented in Section 3 enables inference of the binary’s detector-frame masses and spins. To correct for redshift magnification and obtain a PDF for source-frame parameters, we use the marginal luminosity distance PDF provided by the Bayestar algorithm [63], applied to the highest-SNR trigger. The MBTA SNR optimizer evaluates the joint PDF $p_{m_{1}^{d},m_{2}^{d},\chi_{\mathrm{eff}}}$ at several thousand points across the parameter space. To compensate for the limited number of samples, the conversion problem is simplified by assuming that only the mass ratio is correlated with the spin, resulting in a 2D frame-dependent individual mass PDF:
@@
A change of variables enables expressing the source-frame joint mass and redshift PDF in terms of the detector-frame PDF:
@@
Since our analysis marginalizes over extrinsic parameters, distance is not directly measured. We therefore rely on the luminosity distance PDF $p_{D_{L}}$ estimated by Bayestar, based on the highest-$\mathrm{SNR}$ trigger. To obtain an estimate of the redshift PDF, we again apply a change of variables:
@@
where the relation $d_{L}=f_{z\rightarrow d_{L}}(z)$ is derived from a standard cosmological model, as described in [16]. The distance PDF produced by Bayestar relies on the assumption that the mass parameters are that of the highest-SNR trigger, and would be different for different masses, due to the dependence of the GW inspiral amplitude (primarily) on the chirp mass [27]. To take this into account, we define the conditional redshift PDF for given source-frame component masses as a homothetic transform of the PDF in Equation 28:
@@
with $m_{\mathrm{chirp,highest\text{-}SNR}}^{d}$ the detector-frame chirp mass of the highest-$\mathrm{SNR}$ trigger (for which the redshift PDF is known) and $m_{\text{chirp}}^{s}$ the source-frame chirp mass for which we want to estimate the redshift PDF value.
As a result, the joint distribution of Equation 27 can be factorized and marginalized over redshift to get the source-frame individual mass PDF:
@@
To provide a usable data product, this PDF is sampled using rejection sampling. A comparison of the detector-frame and source-frame posterior probabilities obtained with this method is shown in Figure 3.

### v2

@@ 2.2 Fast template placement
  • projecting this bank using the metric of the template that maximizes the posterior probability density function (PDF) (Equation 9).
@@ 3.1 MBTA detector-frame PE
Under the hypothesis that a CBC signal is present in the data $d$, the posterior PDF for the parameters $\boldsymbol{\lambda}$ of a given template $T$ is given by Bayes theorem, as the product of the template likelihood and prior PDF:
$f(T(\boldsymbol{\lambda})\mid d)\propto\mathcal{L}(d\mid T(\boldsymbol{\lambda}))\cdot f(T(\boldsymbol{\lambda})).$ (9)
@@
$\mathcal{L}(d\mid T({\boldsymbol{\lambda}_{\mathrm{int}}}))\propto\exp\left(-\sum\limits_{i}\mathrm{SNR}_{\mathrm{opt},i}^{2}\left(1-\frac{\mathrm{SNR}_{T,i}}{\mathrm{SNR}_{\mathrm{opt},i}}\right)\right),$ (10)
@@
The second term of Equation 9 is the prior PDF of the source parameters. To ensure consistency with the parameters reported by the collaboration in [6], we adopt the same agnostic priors: uniform distributions in detector-frame individual masses and spin magnitudes, with isotropic spin orientations. This prior is expressed in the dimensionless chirp time space following the procedure described in Appendix B.
To account for the fact that our samples are drawn from the SNR optimizer template bank, and therefore follow a specific distribution in dimensionless chirp times, we apply the reweighting procedure described in [27], resutlting in probability distributions obtained by dividing the sampled PDFs by the template density—either uniform for the first iteration or following Equation 16 for the second iteration. Figure 2 highlights the differences between the template distribution and the prior probability distribution, illustrating that the chosen prior favors symmetric masses and low effective spins, while the bank covers a broader parameter space. Figure 3a shows the detector-frame posterior probability distribution evaluated for the parameters of each template filtered during the second iteration of the MBTA SNR optimizer, applied to the data associated with GW230529_181500 [2].
Figure 2: Prior probability distribution of the templates used in the first iteration of the SNR optimizer, initiated by the MBTA trigger for GW230529_181500 [2], used alongside recovered SNRs to infer source parameters. The distributions are shown for the detector-frame chirp mass ($m_{\text{chirp}}^{d}$), the symmetric mass ratio ($\eta$), and the effective spin ($\chi_{\mathrm{eff}}$). The plots display the marginal 1D (diagonal) and 2D (off-diagonal) distributions. Black lines indicate the distributions of the templates used for filtering. Blue lines and dots show the same distributions reweighted according to Equation 20.
@@ 3.2 MBTA source-frame PE
The source properties can be inferred from mass and spin estimates given by Equation 9. However, because of the Universe’s expansion, the masses measured in the detector-frame appear larger due to redshift $z$ [43]. Since our PE method only probes the intrinsic parameters, it does not infer the source distance, nor the redshift. We therefore rely on the source localization provided in low-latency by the Bayestar algorithm [63], applied to the maximum likelihood template. Following the procedure in Appendix C, we convert the detector-frame mass and spin joint posterior probability distribution into source-frame using the marginal luminosity distance probability distribution provided by Bayestar and assuming a standard cosmological model, described in [16].
@@
(b) Source-frame. Samples are obtained by rejection sampling of the calculated posterior probability distribution.
@@ 3.3 Source classification and properties
The resulting source-frame posterior probability distribution can be used to update the source classification and properties publicly provided to aid for inform EM surveys of the candidates.
Source classification is expressed as a set of probabilities representing the possible origin of the signal. Four mutually exclusive categories are considered: $p_{\text{terr}}$ for a non-astrophysical (terrestrial) origin, $p_{\text{BNS}}$ for both component masses between $1~\mathrm{M_{\odot}}$ and $3~\mathrm{M_{\odot}}$, $p_{\text{NSBH}}$ for a binary with one mass in this range and the other above $3~\mathrm{M_{\odot}}$, and $p_{\text{BBH}}$ for both components above $3~\mathrm{M_{\odot}}$ [46]. Since the MBTA SNR optimizer relies on a significant detection and does not re-estimate candidate significance, the value of $p_{\text{terr}}$ is taken from the online search [20, 19]. By integrating the source-frame joint individual mass posterior probability distribution, marginalized over spin, the remaining probability is redistributed among the three astrophysical categories. Since standard LVK searches do not include a $p_{\text{astro}}$ sub-solar mass (SSM) category, the portion of the posterior probability distribution corresponding to SSM components is excluded from this source classification process.
The collaboration also provides a set of probabilities inferring the possible properties of the source [46]. While the method for estimating these properties is shared across all LVK search pipelines [28], we show that the MBTA SNR optimizer PE process can be used to produce our own estimate. The HasSSM and HasMassGap probabilities can be computed by directly integrating the source-frame mass posterior probability distribution over the regions where at least one component has a mass below $1~\mathrm{M_{\odot}}$, and between $3~\mathrm{M_{\odot}}$ and $5~\mathrm{M_{\odot}}$, respectively. The HasNS and HasRemnant definitions are dependent on neutron star models. Therefore, the limits for neutron star and remnant masses are marginalized over several neutron star equations of state [44] to weight the source-frame posterior probability distribution before integration. In this process, neutron star spins in NSBH systems are fixed to zero, while black hole spins are allowed to vary according to the measured posterior probability distribution.
@@ 4.3 Parameter recovery
For further validation, we applied our PE method to the O4a-events and compared the 2D posterior probability distributions of the source-frame component masses obtained with our method to those provided by the LVK collaboration. We perform this comparison by evaluating the overlap $\mathcal{O}$ between the two binned distributions—denoted $p_{\mathrm{MBTA}}$ and $p_{\mathrm{LVK}}$—given by [25]:
@@ 4.5 Source classification and properties
Releasing the SNR optimizer classifications and properties would be appropriate even in the absence of an $\mathrm{SNR}$ gain, as they remain consistent with the results shown in Figures 10b and 11, the PE procedure—based on posterior probability distribution integration—being robust to small $\mathrm{SNR}$ variations.
@@ Acknowledgments
We express our gratitude to Gijs Nelemans and Jérôme Novak for their work in carrying out the internal LVK review of the MBTA pipeline and for their helpful comments and suggestions on the method presented in this paper. We acknowledge Gregory Ashton for reviewing this paper as part of the LVK internal review process and for providing valuable feedback. We thank our LVK collaborators from the CBC and low-latency groups for constructive comments. This analysis exploits the resources of the computing facility at the EGO-Virgo site, and of the Computing Center of the Institut National de Physique Nucléaire et Physique des Particules (CC-IN2P3 / CNRS). This research has made use of data or software obtained from the Gravitational Wave Open Science Center (gwosc.org), a service of the LIGO Scientific Collaboration, the Virgo Collaboration, and KAGRA. This material is based upon work supported by NSF’s LIGO Laboratory which is a major facility fully funded by the National Science Foundation, as well as the Science and Technology Facilities Council (STFC) of the United Kingdom, the Max-Planck-Society (MPS), and the State of Niedersachsen/Germany for support of the construction of Advanced LIGO and construction and operation of the GEO600 detector. Additional support for Advanced LIGO was provided by the Australian Research Council. Virgo is funded, through the European Gravitational Observatory (EGO), by the French Centre National de Recherche Scientifique (CNRS), the Italian Istituto Nazionale di Fisica Nucleare (INFN) and the Dutch Nikhef, with contributions by institutions from Belgium, Germany, Greece, Hungary, Ireland, Japan, Monaco, Poland, Portugal, Spain. KAGRA is supported by Ministry of Education, Culture, Sports, Science and Technology (MEXT), Japan Society for the Promotion of Science (JSPS) in Japan; National Research Foundation (NRF) and Ministry of Science and ICT (MSIT) in Korea; Academia Sinica (AS) and National Science and Technology Council (NSTC) in Taiwan.
@@ A.2 Universal bank generation
As discussed at the end of Section 2, we consider two different approaches. In the first iteration of the SNR optimizer, a constant minimal match of $98\%$ is used, resulting in templates uniformly distributed in dimensionless chirp times. At this stage, $4000$ potential templates are retained, represented in the top row of Figure 1. In the second iteration, templates are distributed to focus around the most probable point identified by the first pass. We adopt a sampling strategy designed to match the template distribution to the likelihood expressed in Equation 10. Templates are placed such that the probability density of the match $M$ with respect to the bank center, characterized by $\boldsymbol{\theta}_{O}$, is given by:
$f_{T}(\boldsymbol{\theta})=f_{0}\exp\!\left[-\tilde{\rho}^{2}\left(1-M(\boldsymbol{\theta}_{O},\boldsymbol{\theta})\right)\right],$ (16)
@@
with $\bar{\rho}$ denoting the mean SNR measured across detectors for the template yielding the highest network $\mathrm{SNR}$, $k$ a safety factor (typically set to $0.4$) used to artificially broaden the bank, and $\tilde{\rho}_{\max}$ a parameter meant to set an upper limit on $\tilde{\rho}$, fixed to $15$ to ensure that the second iteration bank encompasses more than one template from the first iteration. The normalization constant $f_{0}$ is chosen so that the resulting bank contains $7000$ potential templates.
@@ A.3 Sample projection
Appendix B Prior calculation
In standard LVK parameter inference, an agnostic prior is typically assumed. This prior is taken to be uniform within some plausible range in individual detector-frame component masses $(m_{1}^{d},m_{2}^{d})$ and spin magnitudes $(\chi_{1},\allowbreak\chi_{2})$, while isotropic in spin orientations [5]. However, the parameter estimation technique presented here relies on samples expressed in terms of the dimensionless chirp times $\boldsymbol{\theta}\equiv(\theta_{0},\theta_{3},\theta_{3s})$, defined in Equation 7. Since individual spins are constrained to be identical by the bank generation process, the spin dependency in $\boldsymbol{\theta}$ is encapsulated in a single spin component:
$\chi^{z}=\chi_{1}\cos(\phi_{1})=\chi_{2}\cos(\phi_{2}),$ (19)
with $(\phi_{1},\phi_{2})$ the spin orientation angles relative to the orbital angular momentum of the binary. Therefore the prior PDF used in Equation 9 can be expressed as:
$f(T(\boldsymbol{\lambda}\equiv\boldsymbol{\theta}))=|J|\cdot f_{m_{1}^{d},m_{2}^{d},\chi^{z}}(m_{1}^{d},m_{2}^{d},\chi^{z})$ (20)
@@ Appendix B Prior calculation
with $\omega_{\eta}^{2}=4\pi^{2}f_{0}^{2}\sqrt{1-4\eta}$. Assuming a standard prior where masses and spins are uncorrelated and masses are uniformly distributed, the prior PDF simplifies:
$\begin{split}f_{m_{1}^{d},m_{2}^{d},\chi^{z}}(m_{1}^{d},m_{2}^{d},\chi^{z})\propto f_{\chi^{z}}(\chi^{z}).\end{split}$ (22)
@@
$f_{\chi^{z}}(\chi^{z})=-\frac{1}{2}\log_{e}\left(|\chi^{z}|\right).$ (23)
Combining Equation 21 and Equation 23 gives the prior PDF expressed in dimensionless chirp times. To prevent the prior from diverging at $\eta=0.25$ (Equation 21) or $\chi^{z}=0$ (Equation 23), the parameters used in its evaluation are clipped to $q=m_{2}/m_{1}\leq 0.99$ and $|\chi^{z}|\geq 0.001$. Reweighting by the PDF used for template placement yields the prior probability distribution, illustrated in Figure 2.
@@ Appendix C Detector-frame into source-frame posterior conversion
The parameter estimation method presented in Section 3 enables inference of the binary’s detector-frame masses and spins. To correct for redshift magnification and obtain a distribution of source-frame parameters, we use the marginal luminosity distance probability distribution provided by the Bayestar algorithm [63], applied to the highest-SNR trigger. The MBTA SNR optimizer evaluates the joint probability distribution $p_{m_{1}^{d},m_{2}^{d},\chi_{\mathrm{eff}}}$ at several thousand points across the parameter space. To compensate for the limited number of samples, the conversion problem is simplified by assuming that only the mass ratio is correlated with the spin, resulting in a 2D frame-dependent individual mass probability distribution:
@@
A change of variables enables expressing the source-frame joint mass and redshift probability distribution in terms of the detector-frame probability distribution:
@@
Since our analysis marginalizes over extrinsic parameters, distance is not directly measured. We therefore rely on the luminosity distance probability distribution $p_{D_{L}}$ estimated by Bayestar, based on the highest-$\mathrm{SNR}$ trigger. To obtain an estimate of the redshift probability distribution, we again apply a change of variables:
@@
where the relation $d_{L}=f_{z\rightarrow d_{L}}(z)$ is derived from a standard cosmological model, as described in [16]. The distance probability distribution produced by Bayestar relies on the assumption that the mass parameters are that of the highest-$\mathrm{SNR}$ trigger, and would be different for different masses, due to the dependence of the GW inspiral amplitude (primarily) on the chirp mass [27]. To take this into account, we define the conditional redshift probability for given source-frame component masses as a homothetic transform of the probability distribution in Equation 27:
@@
with $m_{\mathrm{chirp,highest\text{-}SNR}}^{d}$ the detector-frame chirp mass of the highest-$\mathrm{SNR}$ trigger (for which the redshift probability is known) and $m_{\text{chirp}}^{s}$ the source-frame chirp mass for which we want to estimate the redshift probability value.
As a result, the joint distribution of Equation 26 can be factorized and marginalized over redshift to get the source-frame individual mass probability distribution:
@@
To provide a usable data product, this probability distribution is sampled using rejection sampling. A comparison of the detector-frame and source-frame posterior probability distributions obtained with this method is shown in Figure 3.